### v1

# Prethermal gauge structure and surface growth in $\mathbb{Z}_{2}$ lattice gauge theories

Abstract
Universal aspects of thermalization in interacting many-body systems are typically challenging to derive microscopically, yet provide a powerful framework for understanding emergent phenomena. Here, we numerically study the mean-field dynamics of a (2+1)(2+1)D spin system with thousands of spins and show that experimentally-feasible two-body Ising interactions can stabilize a prethermal $\mathbb{Z}_{2}$ lattice gauge structure with dynamical matter, manifested by a gauge-invariant plateau with exponentially long lifetime. Eventually, the metastable prethermal $\mathbb{Z}_{2}$ gauge structure breaks down via a proliferation of Gauss’ law defects, similar to bubble formation in false vacuum decay. In this regime, we discover spatio-temporal correlations described by a non-linear surface growth consistent with the (1+1)(1+1)D Kardar-Parisi-Zhang (KPZ) universality class. We benchmark our results in small systems against semi-classical discrete time Wigner approximation (DTWA) and exact diagonalization (ED), where the breakdown of DTWA signals the emergence of an extensive number of local symmetries that strongly influence the thermalization pathway. Our model provides a testbed for quantum simulators and is directly implementable in large-scale arrays of Rydberg atoms.
Introduction.— How isolated many-body systems reach thermal equilibrium remains a central challenge in non-equilibrium physics Nandkishore and Huse (2015); Eisert et al. (2015); D’Alessio et al. (2016), one that has recently become experimentally accessible through highly controlled quantum simulators Kinoshita et al. (2006); Bernien et al. (2017); yoon Choi et al. (2016); Wei et al. (2022); Zenesini et al. (2024); Andersen et al. (2025); Haghshenas et al. (2025). These experiments allow one to probe universal properties of the dynamics and classify far-from-equilibrium phases of matter through, for example, hydrodynamic descriptions Prüfer et al. (2018); Erne et al. (2018); Johnstone et al. (2019); Navon et al. (2019); Glidden et al. (2021); García-Orozco et al. (2022); Huh et al. (2024); Martirosyan et al. (2024).
In recent years, lattice gauge theories (LGTs), originally developed in the context of high-energy physics, have become a prominent playground to explore non-trivial dynamics, due to their enriched structure rooted in local constraints Bañuls et al. (2020); Halimeh et al. (2025a, b); Dalmonte and Montangero (2016). Prominent phenomena observed in far-from-equilibrium experiments include quantum many-body scarring Bernien et al. (2017); Surace et al. (2020); Su et al. (2023), many-body localization Smith et al. (2017); Brenes et al. (2018); Gyawali et al. (2024), and string breaking Martinez et al. (2016); González-Cuadra et al. (2025); Cochran et al. (2025); De et al. (2024). This makes LGTs a versatile framework to study non-equilibrium phenomena, some of which have also been observed in a broader class of strongly-interacting many-body systems, from Hubbard models Schreiber et al. (2015); Adler et al. (2024) to spin systems Bernien et al. (2017); Ljubotina et al. (2019); Wei et al. (2022) and superfluids Zenesini et al. (2024); Zhu et al. (2024).
Despite these challenges, a universal description of thermalization in LGTs is lacking and largely unexplored – particularly for kinetically constrained models Gromov et al. (2020). This would be of importance to a variety of fields. Understanding the thermalization of non-Abelian gauge theories plays a crucial role in heavy-ion colliders to extrapolate between the observed thermal state and the strongly-correlated quark-gluon plasma produced during the collision Berges et al. (2021). Strongly-correlated quantum matter, such as quantum spin liquids, can exhibit unusual hydrodynamics Bulchandani et al. (2021); Qi et al. (2023). In the context of quantum simulation, where gauge symmetry must be artificially enforced Wiese (2013); Halimeh et al. (2025a), characterizing the system’s dynamical response to gauge defects is essential to achieving scalability to larger system sizes Halimeh and Hauke (2020).
Here, we employ efficient classical spin simulations of a $\mathbb{Z}_{2}$ LGT with dynamical matter over large space and time scales to unveil the universal features of its thermalization. The model is motivated by its experimental feasibility in Rydberg arrays Homeier et al. (2023), and constructed such that the gauge-invariant subspace arises as a metastable manifold, protected by an energy gap $\propto V$. In the presence of a weak gauge-symmetry-breaking drive $\Omega$, the system exhibits a long prethermal gauge-invariant plateau, which emerges above a critical protection strength $V_{c}$, relevant to the large-scale quantum simulation of gauge theories Halimeh et al. (2025a). At late times, thermalization consists of the breakdown of gauge symmetry, which we show to be governed by the nucleation and growth of gauge defects. A careful scaling analysis of the surface growth shows critical exponents consistent with the (1+1)(1+1)D KPZ universality class Kardar et al. (1986); Corwin (2012). We further benchmark small systems beyond mean-field dynamics using DTWA Schachenmayer et al. (2015) and ED. Surprisingly, we find that, while the mean-field dynamics shows qualitative agreement with ED, DTWA is not able to capture the prethermal plateau, highlighting how, unlike in common cases, thermalization in LGTs cannot be understood as a simple phenomenon of scrambling of fluctuations, but is instead strongly affected by the local (emergent) dynamical constraints.
Model.—We consider spins $\hat{\bm{\sigma}}=(\hat{\sigma}^{x},\hat{\sigma}^{y},\hat{\sigma}^{z})$ on the sites $j=(X,Y)$ and links $\langle i,j\rangle$ of a honeycomb lattice, see Fig. 1(a,b). By analogy with Ising LGTs Wegner (1971); Fradkin and Susskind (1978); Fradkin and Shenker (1979), the matter sites can be occupied by $\mathbb{Z}_{2}$ charges ($\hat{\sigma}^{z}_{j}=+1$) and the links host $\mathbb{Z}_{2}$ electric fields ($\hat{\sigma}^{z}_{\langle i,j\rangle}=-1$), for which we define the $\mathbb{Z}_{2}$ Gauss’ law on vertex $j$
$\hat{G}_{j}=-\hat{\sigma}^{z}_{j}\prod_{i:\langle i,j\rangle}\hat{\sigma}^{z}_{\langle i,j\rangle}$ (1)
containing the matter site $j$ and its adjacent links, such that $G_{j}:=\langle\hat{G}_{j}\rangle\in[-1,1]$; a vertex $j$ is fully gauge symmetric (symmetry breaking) if $G_{j}=+1$ ($G_{j}=-1$). A Hamiltonian has a local $\mathbb{Z}_{2}$ symmetry if $\hat{H},\hat{G}_{j}]=0$ for all $j$, such that the Gauss’ laws are a constant of motion, $\dot{G_{j}}=0$.
The system has a Hamiltonian $\hat{H}=\hat{H}_{\mathbb{Z}_{2}}+\hat{H}_{V}+\hat{H}_{\Omega}$, with
$\begin{split}\hat{H}_{\mathbb{Z}_{2}}&=J\sum_{\langle i,j\rangle}\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{\langle i,j\rangle}\hat{\sigma}^{x}_{j},\\ \hat{H}_{V}&=V\sum_{j}\frac{(1+\delta_{j})}{8}\left[\left(\hat{\sigma}^{z}_{j}+\sum_{i:\langle i,j\rangle}\hat{\sigma}^{z}_{\langle i,j\rangle}\right)^{2}-4\right],\\ \hat{H}_{\Omega}&=\frac{\Omega}{2}\left(\sum_{j}\hat{\sigma}^{x}_{j}+\sum_{\langle i,j\rangle}\hat{\sigma}^{x}_{\langle i,j\rangle}\right).\end{split}$ (2)
The first term $\hat{H}_{\mathbb{Z}_{2}}$ is a gauge-invariant coupling between matter and gauge fields, typical for Ising $\mathbb{Z}_{2}$ LGTs Fradkin and Susskind (1978); Fradkin and Shenker (1979). The second term $\hat{H}_{V}$ describes a local pseudogenerator Homeier et al. (2023); Halimeh et al. (2022a), composed of two-body interactions in the $z$ basis. Up to an overall energy shift, this term is proportional to $\sum_{j}(\hat{\sigma}^{z}_{j,\rm tot})^{2}$, where $\hat{\sigma}^{z}_{j,\rm tot}$ is the total magnetization of all spins within vertex $j$. For $\delta_{j}=0$, the term engineers an energy landscape where the gauge symmetric vertices contribute no energy, and the gauge-symmetry breaking vertices contribute an energy 0.5V-0.5V or $1.5V$, see Fig. 1(a). That is, $\hat{H}_{V}$ introduces a gauge protection via energy penalties Halimeh et al. (2021). The third term $\hat{H}_{\Omega}$ describes rotations about the $x$ axis, breaking the gauge symmetry, see Fig. 1(b).
In the following, we will mostly focus on the mean-field dynamics $\dot{\vec{\sigma}}=\{\vec{\sigma},H\}$, where $\vec{\sigma}$ denotes a classical unit vector satisfying Poisson brackets $\{\sigma^{a},\sigma^{b}\}=\varepsilon_{abc}\sigma^{c}$, with $\varepsilon_{abc}$ the Levi-Civita symbol. Conveniently, the classical mean-field dynamics can be efficiently solved for large (2+1)(2+1)D systems. In the last part, we will compare it to semi-classical DTWA and ED in small systems.
Figure 1: Spin model and emergent gauge structure. (a) Spins are located on the sites (matter) and links (electric/gauge fields) of a honeycomb lattice. We define a $\mathbb{Z}_{2}$ Gauss’ law, see Eq. (1), together with a two-body Hamiltonian $\hat{H}_{V}$ that energetically protects a metastable sector at energy $E=0$ containing the gauge-invariant configurations $G_{j}=+1$ (inset). (b), (c) The mean-field dynamics describe the evolution of classical spins, i.e., unit vectors on a sphere. After the system is prepared in a random gauge-invariant state, the perturbation $\hat{H}_{\Omega}$ rotates the spins about the $x$ axis, leading to the dynamical violation of the local Gauss’ law constraints at later times $t$. (d) We evolve gauge-invariant initial states for various protections strengths $V/\Omega=4,6,10,20,40,80$ (light to dark color) averaging over 200200 random realizations in a $20\times 20$ plaquette system. The gauge-symmetry breaking errors remain small and controlled for a long time, leading to prethermal plateau with an emergent $\mathbb{Z}_{2}$ gauge structure. (e) The protection strength $V/\Omega$ controls the magnitude (top) and critical time (bottom) of the prethermal plateau; black lines indicate fit functions. (f) Local Gauss’ laws of one exemplary trajectory for multiple times after the plateau phase, see yellow markers in panel (d). Local nucleation regions followed by a proliferation of Gauss’ law defects imply a rich spatio-temporal correlation structure.
Prethermal $\mathbb{Z}_{2}$ LGT.— Initializing the system in a configuration with $G_{j}=+1$ for all $j$, we expect the system to remain gauge-invariant for a long time if $\Omega\ll V$. For $\delta_{j}=0$, the breaking of the Gauss’ law constraints will ultimately happen through correlated processes on four vertices due to resonant processes where one vertex increases energy by +1.5V+1.5V and three vertices decrease their energy by 0.5V-0.5V, see Fig. 1(a). The duration of the gauge invariance can be further extended lifting these degeneracies with random couplings $\delta_{j}$, e.g., normally distributed around zero with standard deviation $\delta=0.1$} Homeier et al. (2023); see the End Matter for a discussion of various disorder strengths.
We probe the breakdown of the gauge symmetry by numerically integrating the mean-field dynamics on a $20\times 20$ plaquette system ($N=2000$ spins, $N_{\rm vert}=800$ vertices and $3N$ coupled differential equations) with periodic boundaries and using trotterization techniques Howell et al. (2019); Pizzi et al. (2021), see End Matter for details. We consider an ensemble of $200$ trajectories, each with a different random gauge-invariant initial configuration and set of couplings $\delta_{j}$. The gauge-symmetry breaking is quantified by the time-averaged error $\varepsilon(t)=t^{-1}\int_{0}^{t}\mathrm{d}\tau[1-\langle G_{j}(\tau)\rangle]$, where $\langle\cdot\rangle$ denotes the average over the ensemble and vertices, see Fig. 1(c). Its behavior is shown in Fig. 1(d) for various protection strengths $V/\Omega=4,\ldots,80$; in the following, we set $J=\Omega$.
We find three distinct regimes Halimeh and Hauke (2020): (i) At early times $t\Omega\lesssim 1$}, the error shows a universal growth $\varepsilon(t)\propto t^{2}$. (ii) The error settles to a stable plateau with small gauge-symmetry breaking for an exponentially long time in $V/\Omega$. Hence, the system establishes a prethermal phase with an emergent $\mathbb{Z}_{2}$ gauge symmetry. In Fig. 1(e) we plot the value $\varepsilon_{\mathrm{pre}}:=\varepsilon(t\Omega=20)$ of the prethermal plateau and find that it follows a power-law $\varepsilon_{\mathrm{pre}}\propto(\Omega/V)^{2}$ controlled by the protection strength $V$ Halimeh et al. (2022a). The critical time after which the plateau becomes unstable is defined as the time of maximum curvature $t_{\mathrm{crit}}=\mathrm{max}_{t}\partial^{2}\varepsilon(t)/\partial t^{2}$ after the plateau, and scales as $t_{\mathrm{crit}}\propto\exp(V/\Omega)$. (iii) At late times, the $\mathbb{Z}_{2}$ Gauss’ law breaks down and the system reaches the infinite temperature value, $\varepsilon\rightarrow 1$. Finally, an examplary trajectory in Fig. 1(f) shows Gauss’ law violations with strong spatio-temporal correlations, which we will discuss in detail below.
The existence of a prethermal phase with small and controlled gauge-symmetry breaking errors is crucial for the gauge-invariant quantum simulation of $\mathbb{Z}_{2}$ LGTs. Going beyond the small-scale ED predictions Homeier et al. (2023); Halimeh et al. (2022a, b), our results show a promising method to stabilize gauge constraints in large-scale (2+1)(2+1)D quantum simulators Homeier et al. (2023).
Figure 2: Mean-field phase transition. (a) Potential $F(G)$ for the translationally-invariant model without dynamical matter, for protection strengths $V/\Omega=2,..,5$ (bright to dark) and $(V/\Omega)_{c}\approx 3.33$ (orange). The dynamics is governed by the kinetic equation $(\dot{G})^{2}+F(G)=0$ with initial conditions $G(0)=+1$ and $\dot{G}(0)=0$. (b) For a protection strength below the critical value, the error grows to its maximum in the translationally-invariant setting (red curve). As the protection is increased, the error is suppressed since the root in the potential $F(G)$ moves towards $G=1$. Inset: The Gauss’ law shows large gauge-symmetry breaking oscillations for $(V/\Omega)=3.3<(V/\Omega)_{c}$ (red curve), and constrained dynamics for $(V/\Omega)=3.4>(V/\Omega)_{c}$ (light gray curve) with smaller oscillations for increasing protection, e.g., $V/\Omega=5$ (dark gray curve).
Critical protection.— Next, we gain analytic insights into the thermalization behavior in a simple setting, which allows us to predict a phase transition, where above a critical protection strength $(V/\Omega)_{c}$ the dynamics exhibits a prethermal plateau. To this end, we consider the translationally-invariant model without dynamical matter ($\sigma^{z}_{j}(t)\equiv-1$ for all $t$ and $J=0$), and we derive the mean-field equations of motion for the initial state with no electric field lines ($\sigma^{z}_{\langle i,j\rangle}(0)=+1$). Conservation of energy and spin length yields the kinematic equation for Gauss’ law $\dot{G}^{2}+F(G)=0$, where $F(G)=F(G;\Omega,V)$ is a potential energy, see End Matter for a derivation. In Fig. 2(a), we plot the potential for various protection strengths $V/\Omega$: Starting from a gauge-invariant state $G=+1$, the gauge-breaking dynamics is unconstrained when $V/\Omega$ is below a critical protection strength $(V/\Omega)_{c}=\sqrt{\frac{11+5\sqrt{5}}{2}}\approx 3.33$, and the Gauss’ law shows strong oscillations, see Fig. 2(b). Above $(V/\Omega)_{c}$, the dynamics becomes increasingly constrained and oscillates around a value close to $G=+1$, which is associated with the development of an additional root in the potential $F(G)$ shown by the orange curve in Fig. 2(a).
Figure 3: Surface growth. (a) We investigate the dynamics starting from an initial configuration with Gauss’ law defects in the bottom row. We find defect and defect-free regions separated by a surface (red line, right) that we quantify by the height function $h_{n}(X,t)$, where $n$ labels a trajectory for a set of random protection strengths $\{\delta_{j}\}$; here we set $V/\Omega=10$. (b) We analyze the mean height $h$ and width $\delta h$, Eq. (3), showing ballistic spreading and a non-linear roughening, respectively; statistical error bars of the numerical data are smaller than the solid lines. The gray lines show power-law fits to the data for $50<t\Omega<1000$, where the width $\delta h$ does not yet crossover to the asymptotic Family-Vicsek behavior. (c) Top: Width $\delta h$ for increasing linear system size $L=40,80,160,320$ (light to dark). Bottom: The dynamical scaling analysis reveals an excellent collapse of the data for a roughening exponent $\alpha=1/2$ and dynamical exponent $z=3/2$.
Surface growth.—The trajectory of an initial gauge-invariant configuration reveals a characteristic behavior, see Fig. 1(f): Upon exiting the pre-thermal plateau, random fluctuations, determined by the initial configuration and the couplings $\delta_{j}$, lead to the nucleation of gauge-symmetry breaking errors. Subsequently, neighboring vertices violate the Gauss’ law constraint, so that the errors proliferate until the entire system reaches equilibrium, suggesting a universal hydrodynamic description based on a surface growth model Corwin (2016).
To test our hypothesis, we perform numerical simulation on systems with $50\times 50$ plaquettes and periodic (open) boundaries in the $X$ ($Y$) spatial direction. We initialize a configuration with a line defect in the bottom row, see Fig. 3(a) left, and repeat the simulations for $N_{\rm Ens}=200$ realizations of random couplings $\delta_{j}$ Foo (a). To each space-time point we associate a binary variable $\Theta(1-G_{j}(t)-g_{\rm{thres}})$, where $\Theta(\cdot)$ is the Heaviside function; in the following we set $g_{\rm{thres}}=0.2$ but our findings do not depend on this specific value. This binary variable defines a surface separating the regions of small and large Gauss’ law violations, see Fig. 3(a). Further, we use the surface to define a height function $h_{n}(X,t)$ in Euclidean space Foo (b) and for each realization $n=1,\ldots,N_{\rm Ens}$.
Next, we analyze the height function and the cumulants associated with this distribution, such as the mean $h(t)$ and width $\delta h(t)$ given by
$\begin{split}h(t)&=\big\langle\langle{h}_{n}(X,t)\rangle_{X}\big\rangle_{n},\\ \delta h(t)&=\bigg\langle\sqrt{\big\langle\left[h_{n}(X,t)-\langle h_{n}(X,t)\rangle_{X}\right]^{2}\big\rangle_{X}}\bigg\rangle_{n},\end{split}$ (3)
where $\langle\cdot\rangle_{n}$ and $\langle\cdot\rangle_{X}$ denote averaging over the ensemble and over the dimension $X$, respectively; for the case of $320$ plaquettes the linear dimension is $L=640$.
In Fig. 3(b), we fit power-law functions $\propto a+bt^{c}$ to the mean height $h(t)$ and width $\delta h(t)$ for times $50<t\Omega<1000$ (grey curves). Such an intermediate time range limits the non-universal behavior at short times and finite-size effects at long times. We observe a ballistic spreading of the mean height $h(t)\propto t^{1.0146(9)}$ and non-linear roughening of the surface width $\delta h(t)\propto t^{\,\beta}$ with growth exponent $\beta=0.33(1)$. The error bars describe one standard deviation of a least square minimization fitting procedure, and our simulations are likely dominated by systematics, such as finite size effects, more than by statistical fluctuations. To confirm that our model is consistent with ballistic deposition Corwin (2016) and KPZ scaling, we carefully take finite size effects into account performing a dynamical scaling analysis.
For this, we consider the Family-Vicsek scaling relation $\delta h\sim L^{\alpha}f(t/L^{z})$, where $\alpha$ is the roughening exponent and $z=\alpha/\beta$ is the dynamic exponent Family and Vicsek (1985). The dynamical scaling captures the spreading of spatial correlations, which saturate after a time $t\sim L^{z}$. In Fig. 3(c) top, we show the width $\delta h$ for increasing system sizes $L=40,80,160,320$. By rescaling with the Family-Vicsek relation, we obtain a collapse of the data using the exponents $\alpha=1/2$ and $\beta=1/3$ of the (1+1)(1+1)D KPZ universality class Kardar et al. (1986); Corwin (2016), see Fig. 3(c) bottom.
We note that a complete characterization of the universality class governing the surface growth necessitates to evaluate higher-order cumulants Ljubotina et al. (2019); Wei et al. (2022); Rosenberg et al. (2024), and that in our simulations the skewness and kurtosis of the height field are not converged. Despite these challenges, that future studies should address considering larger systems, our findings strongly suggest KPZ universality of the spreading of Gauss’ law defects.
An alternative interpretation to the proliferation of Gauss’ law defects is provided by introducing a generalized Gauss’ law $\hat{\tilde{G}}_{j}$ under which all terms in the Hamiltonian become gauge invariant, i.e., $\hat{H},\hat{\tilde{G}}_{j}]=0$ and the term $\hat{H}_{\Omega}$ corresponds to pair creation processes mediated by single-body operators. In this way, our results can be directly connected to false-vacuum decay Zhu et al. (2024) and the thermalization dynamics of LGTs with dynamical matter Zhou et al. (2022); Mueller et al. (2025).
[…]
(Semi)classical and quantum model.—The mean-field dynamics considered thus far is exact in the large-spin $S\rightarrow\infty$ limit Mori (2018). We expect that a similar KPZ universality in the thermalization dynamics should appear for a spin-$1/2$ $\mathbb{Z}_{2}$ LGT. This could be probed experimentally in state-of-the-art Rydberg tweezer experiments Browaeys and Lahaye (2020) with hundreds or thousands of qubits Manetsch et al. (2025); Gyger et al. (2024); Chiu et al. (2025). Here, given the notorious numerical challenges of quantum many-body systems Bañuls (2023), we study the long-time dynamics in small systems of $L=2\times 2$ plaquettes with periodic boundary conditions, comparing mean-field dynamics, DTWA, and ED.
In particular, we focus on the experimentally relevant case of only two-body interactions Foo (c), setting $J=0$, and we introduce a yet simpler model without dynamical matter ($\hat{\sigma}^{z}_{j}\equiv 1$} for all $j$). In this model spins only reside on the links of the honeycomb lattice, i.e., on the sites of the dual Kagome lattice, yielding the Hamiltonian
$\hat{H}^{\rm Ryd}=\frac{V}{4}\sum_{\langle A,B\rangle}\hat{\sigma}^{z}_{A}\hat{\sigma}^{z}_{B}+\frac{V}{2}\sum_{A}\hat{\sigma}^{z}_{A}+\frac{\Omega}{2}\sum_{A}\hat{\sigma}^{x}_{A},$ (4)
[…]
In Fig. 4(b,c), we compare the error $\varepsilon(t)$ predicted by the three methods. While the system sizes are too small to analyze the surface growth, we find a prethermal $\mathbb{Z}_{2}$-invariant plateau for both classical dynamics and quantum ED. Notably, the semi-classical DTWA fails to describe the quantum case and is outperformed by the mean-field dynamics. We attribute this behavior to the dynamical emergence of symmetry constraints imposed by the protection term $\hat{H}_{V}$. A potential cause of the DTWA breakdown is that the added stochastic noise, which reproduces the full density matrix of the initial state Schachenmayer et al. (2015), can seed the occupation of gauge-symmetry breaking sectors Muleady et al. (2023) due to the inappropriate evolution of the higher-order correlators. This failure highlights that pre-thermalization in gauge theories strongly depends on the subtle interplay between scrambling of fluctuations and locally constrained dynamics.
Discussion and Outlook.—We have investigated the long-time mean-field dynamics of large-scale spin systems and uncovered a prethermal plateau with emergent $\mathbb{Z}_{2}$ gauge symmetry, whose breakdown follows a universal surface growth of nucleation regions characterized by non-linear growth exponents – opening a path to explore KPZ dynamics in quantum simulators Wei et al. (2022); Rosenberg et al. (2024). These findings are crucial for understanding gauge violation errors and for developing universal frameworks for the thermalization of LGTs Halimeh et al. (2025a). While ED confirms the plateau’s stability in small systems, existing numerical methods, including DTWA, cannot capture the quantum evolution, highlighting quantum simulators Altman et al. (2021) – such as Rydberg tweezer arrays, trapped ions, or superconducting qubits – as promising platforms to experimentally probe these many-body dynamics.
Acknowledgements.—We thank Rahul Nandkishore, Alexander Baumgärtner, Adam Kaufman, Christian Groß, Joanna Lis, Hannes Pichler, Mu Qiao and Ruben Verresen for fruitful discussions. LH and AMR acknowledge support by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (651440), and the NSF JILA-PFC PHY-2317149. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programm (Grant Agreement no 948141) — ERC Starting Grant SimUcQuam. LH and FG were funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2111 – 390814868. AP acknowledges support by Trinity College Cambridge. HZ was funded by the Innovation Program for Quantum Science and Technology (No. 2024ZD0301800). J.C.H. acknowledges funding by the Max Planck Society, the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2111 – 390814868, and the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (Grant Agreement No. 101165667) – ERC Starting Grant QuSiGauge. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@
 Abstract
-Universal aspects of thermalization in interacting many-body systems are typically challenging to derive microscopically, yet provide a powerful framework for understanding emergent phenomena.
-Here, we numerically study the mean-field dynamics of a (2+1)(2+1)D spin system with thousands of spins and show that experimentally-feasible two-body Ising interactions can stabilize a prethermal $\mathbb{Z}_{2}$ lattice gauge structure with dynamical matter, manifested by a gauge-invariant plateau with exponentially long lifetime.
+Universal aspects of thermalization in interacting many-body systems are challenging to derive microscopically, especially in kinetically constrained models, yet their numerical study beyond (1+1)(1+1)D remains notoriously difficult.
+Here, we numerically study the mean-field dynamics of a (2+1)(2+1)D spin system with thousands of spins and show that experimentally-feasible two-body Ising interactions can stabilize a prethermal $\mathbb{Z}_{2}$ lattice gauge structure with dynamical matter, manifested by a separation of timescales with a stable gauge-invariant plateau.
 Eventually, the metastable prethermal $\mathbb{Z}_{2}$ gauge structure breaks down via a proliferation of Gauss’ law defects, similar to bubble formation in false vacuum decay.
-In this regime, we discover spatio-temporal correlations described by a non-linear surface growth consistent with the (1+1)(1+1)D Kardar-Parisi-Zhang (KPZ) universality class.
+In this regime, we discover spatio-temporal correlations described by a non-linear surface growth consistent with the (1+1)(1+1)D Kardar-Parisi-Zhang (KPZ) universality class, revealing a previously hidden feature in the thermalization of multi-point correlators.
 We benchmark our results in small systems against semi-classical discrete time Wigner approximation (DTWA) and exact diagonalization (ED), where the breakdown of DTWA signals the emergence of an extensive number of local symmetries that strongly influence the thermalization pathway.
@@
 (2020); Halimeh et al.
-(2025a, b); Dalmonte and Montangero (2016).
-Prominent phenomena observed in far-from-equilibrium experiments include quantum many-body scarring Bernien et al.
+(2025a, b); Dalmonte and Montangero (2016), with prominent examples observed in far-from-equilibrium experiments Bernien et al.
 (2017); Surace et al.
 (2020); Su et al.
-(2023), many-body localization Smith et al.
+(2023); Smith et al.
 (2017); Brenes et al.
 (2018); Gyawali et al.
-(2024), and string breaking Martinez et al.
+(2024); Smith et al.
+(2017); Brenes et al.
+(2018); Gyawali et al.
+(2024); Martinez et al.
 (2016); González-Cuadra et al.
@@
 (2024).
-Despite these challenges, a universal description of thermalization in LGTs is lacking and largely unexplored – particularly for kinetically constrained models Gromov et al.
+Despite these challenges, a universal description of thermalization in LGTs is largely unexplored – particularly for kinetically constrained models Gromov et al.
+(2020); Chakraborty et al.
 (2025).
-This would be of importance to a variety of fields.
-Understanding the thermalization of non-Abelian gauge theories plays a crucial role in heavy-ion colliders to extrapolate between the observed thermal state and the strongly-correlated quark-gluon plasma produced during the collision Berges et al.
-(2021).
-Strongly-correlated quantum matter, such as quantum spin liquids, can exhibit unusual hydrodynamics Bulchandani et al.
-(2021); Qi et al.
-(2023).
-In the context of quantum simulation, where gauge symmetry must be artificially enforced Wiese (2013); Halimeh et al.
-(2025a), characterizing the system’s dynamical response to gauge defects is essential to achieving scalability to larger system sizes Halimeh and Hauke (2020).
+As the field matures, significant effort has been devoted to find physically relevant phenomena accessible in simple experimental implementations of LGTs Halimeh et al.
+(2025b).
+However, theoretical guidance and concrete predictions are notoriously difficult, since numerical simulations of $(2+1)$ LGTs are currently limited to short times or small system sizes Bañuls (2023).
+Overcoming these limitations requires new approximate approaches that provide access to large system sizes and long times, and that can be benchmarked and tested in quantum simulation experiments.
 Here, we employ efficient classical spin simulations of a $\mathbb{Z}_{2}$ LGT with dynamical matter over large space and time scales to unveil the universal features of its thermalization.
@@
 A careful scaling analysis of the surface growth shows critical exponents consistent with the (1+1)(1+1)D KPZ universality class Kardar et al.
-(1986); Corwin (2012).
+(1986); Corwin (2012); a hidden feature in the spreading of multi-point correlation functions.
 We further benchmark small systems beyond mean-field dynamics using DTWA Schachenmayer et al.
@@
 A Hamiltonian has a local $\mathbb{Z}_{2}$ symmetry if $\hat{H},\hat{G}_{j}]=0$ for all $j$, such that the Gauss’ laws are a constant of motion, $\dot{G_{j}}=0$.
-The system has a Hamiltonian $\hat{H}=\hat{H}_{\mathbb{Z}_{2}}+\hat{H}_{V}+\hat{H}_{\Omega}$, with
+Motivated by its experimental relevance Homeier et al. (2023), we consider a Hamiltonian $\hat{H}=\hat{H}_{\mathbb{Z}_{2}}+\hat{H}_{V}+\hat{H}_{\Omega}$, with
 $\begin{split}\hat{H}_{\mathbb{Z}_{2}}&=J\sum_{\langle i,j\rangle}\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{\langle i,j\rangle}\hat{\sigma}^{x}_{j},\\ \hat{H}_{V}&=V\sum_{j}\frac{(1+\delta_{j})}{8}\left[\left(\hat{\sigma}^{z}_{j}+\sum_{i:\langle i,j\rangle}\hat{\sigma}^{z}_{\langle i,j\rangle}\right)^{2}-4\right],\\ \hat{H}_{\Omega}&=\frac{\Omega}{2}\left(\sum_{j}\hat{\sigma}^{x}_{j}+\sum_{\langle i,j\rangle}\hat{\sigma}^{x}_{\langle i,j\rangle}\right).\end{split}$ (2)
@@
 The gauge-symmetry breaking errors remain small and controlled for a long time, leading to prethermal plateau with an emergent $\mathbb{Z}_{2}$ gauge structure.
+The critical time $t_{\rm crit}$ of the plateau is defined by the maximum curvature (blue dots).
 (e) The protection strength $V/\Omega$ controls the magnitude (top) and critical time (bottom) of the prethermal plateau; black lines indicate fit functions.
@@
 We find three distinct regimes Halimeh and Hauke (2020): (i) At early times $t\Omega\lesssim 1$}, the error shows a universal growth $\varepsilon(t)\propto t^{2}$.
-(ii) The error settles to a stable plateau with small gauge-symmetry breaking for an exponentially long time in $V/\Omega$.
+(ii) The error settles to a stable plateau with small gauge-symmetry breaking for a time linear in $V/\Omega$.
 Hence, the system establishes a prethermal phase with an emergent $\mathbb{Z}_{2}$ gauge symmetry.
@@
 (2022a).
-The critical time after which the plateau becomes unstable is defined as the time of maximum curvature $t_{\mathrm{crit}}=\mathrm{max}_{t}\partial^{2}\varepsilon(t)/\partial t^{2}$ after the plateau, and scales as $t_{\mathrm{crit}}\propto\exp(V/\Omega)$.
+The critical time after which the plateau becomes unstable is defined as the time of maximum curvature after the plateau, see blue dots in Fig.
+1(d), and scales as $t_{\mathrm{crit}}\propto V/\Omega$.
 (iii) At late times, the $\mathbb{Z}_{2}$ Gauss’ law breaks down and the system reaches the infinite temperature value, $\varepsilon\rightarrow 1$.
@@
 The existence of a prethermal phase with small and controlled gauge-symmetry breaking errors is crucial for the gauge-invariant quantum simulation of $\mathbb{Z}_{2}$ LGTs.
-Going beyond the small-scale ED predictions Homeier et al.
+Going beyond (1+1)(1+1)D systems Van Damme et al.
+(2025) and small-scale ED predictions Homeier et al.
 (2023); Halimeh et al.
@@
 Inset: The Gauss’ law shows large gauge-symmetry breaking oscillations for $(V/\Omega)=3.3<(V/\Omega)_{c}$ (red curve), and constrained dynamics for $(V/\Omega)=3.4>(V/\Omega)_{c}$ (light gray curve) with smaller oscillations for increasing protection, e.g., $V/\Omega=5$ (dark gray curve).
-Critical protection.— Next, we gain analytic insights into the thermalization behavior in a simple setting, which allows us to predict a phase transition, where above a critical protection strength $(V/\Omega)_{c}$ the dynamics exhibits a prethermal plateau.
+Critical protection.— Next, we gain analytic insights and physical intuition into the prethermalization behavior by studying the simplest possible setting.
+This allows us to predict a phase transition, where above a critical protection strength $(V/\Omega)_{c}$ the dynamics exhibits a prethermal plateau.
 To this end, we consider the translationally-invariant model without dynamical matter ($\sigma^{z}_{j}(t)\equiv-1$ for all $t$ and $J=0$), and we derive the mean-field equations of motion for the initial state with no electric field lines ($\sigma^{z}_{\langle i,j\rangle}(0)=+1$).
@@
 Surface growth.—The trajectory of an initial gauge-invariant configuration reveals a characteristic behavior, see Fig.
-1(f): Upon exiting the pre-thermal plateau, random fluctuations, determined by the initial configuration and the couplings $\delta_{j}$, lead to the nucleation of gauge-symmetry breaking errors.
+1(f): Upon exiting the prethermal plateau, random fluctuations, determined by the initial configuration and the couplings $\delta_{j}$, lead to the nucleation of gauge-symmetry breaking errors.
 Subsequently, neighboring vertices violate the Gauss’ law constraint, so that the errors proliferate until the entire system reaches equilibrium, suggesting a universal hydrodynamic description based on a surface growth model Corwin (2016).
+The role of spatio-temporal correlations have been previously discussed in kinetically constrained models in (1+1)(1+1)Michailidis et al.
+(2018); Causer et al.
+(2024); Marić et al.
+(2026).
 To test our hypothesis, we perform numerical simulation on systems with $50\times 50$ plaquettes and periodic (open) boundaries in the $X$ ($Y$) spatial direction.
@@
 Despite these challenges, that future studies should address considering larger systems, our findings strongly suggest KPZ universality of the spreading of Gauss’ law defects.
+While spatio-temporal correlations remain hidden in multi-point correlation functions, they are naturally revealed within our gauge-theoretical interpretation of the transverse- and longitudinal-field Ising model on the honeycomb lattice.
 An alternative interpretation to the proliferation of Gauss’ law defects is provided by introducing a generalized Gauss’ law $\hat{\tilde{G}}_{j}$ under which all terms in the Hamiltonian become gauge invariant, i.e., $\hat{H},\hat{\tilde{G}}_{j}]=0$ and the term $\hat{H}_{\Omega}$ corresponds to pair creation processes mediated by single-body operators.
@@
 Here, given the notorious numerical challenges of quantum many-body systems Bañuls (2023), we study the long-time dynamics in small systems of $L=2\times 2$ plaquettes with periodic boundary conditions, comparing mean-field dynamics, DTWA, and ED.
-In particular, we focus on the experimentally relevant case of only two-body interactions Foo (c), setting $J=0$, and we introduce a yet simpler model without dynamical matter ($\hat{\sigma}^{z}_{j}\equiv 1$} for all $j$).
-In this model spins only reside on the links of the honeycomb lattice, i.e., on the sites of the dual Kagome lattice, yielding the Hamiltonian
-$\hat{H}^{\rm Ryd}=\frac{V}{4}\sum_{\langle A,B\rangle}\hat{\sigma}^{z}_{A}\hat{\sigma}^{z}_{B}+\frac{V}{2}\sum_{A}\hat{\sigma}^{z}_{A}+\frac{\Omega}{2}\sum_{A}\hat{\sigma}^{x}_{A},$ (4)
-[…]
-In Fig.
-4(b,c), we compare the error $\varepsilon(t)$ predicted by the three methods.
+In particular, we focus on the experimentally relevant case of only two-body interactions Foo (c) ($J=0$), no disorder ($\delta_{j}=0$), and we set $V/\Omega=10$.
+Analogous to above, we compute the time and ensemble averaged error $\varepsilon(t)$ predicted by the three methods in a model with matter, see Fig.
+4(a), and without matter (setting $\hat{\sigma}^{z}_{j}\equiv 1$} for all $j$), see Fig.
+4(b).
 While the system sizes are too small to analyze the surface growth, we find a prethermal $\mathbb{Z}_{2}$-invariant plateau for both classical dynamics and quantum ED.
+The remarkable, qualitative agreement in our case is possible because the Abelian $\mathbb{Z}_{2}$ Gauss’ law is a non-connected correlator without genuine quantum correlations.
 Notably, the semi-classical DTWA fails to describe the quantum case and is outperformed by the mean-field dynamics.
@@
 This failure highlights that pre-thermalization in gauge theories strongly depends on the subtle interplay between scrambling of fluctuations and locally constrained dynamics.
+Lastly, we describe how the model studied in Fig.
+4(b) can be implemented in quantum simulators–we highlight that this particular regime presents the simplest instance to realize the kinetic constraints.
+Since the matter spins on the sites are frozen, we remove them to obtain a model with spins only residing on the links of the honeycomb lattice, i.e., on the sites of the dual Kagome lattice, yielding the Hamiltonian
+$\hat{H}^{\rm Ryd}=\frac{V}{4}\sum_{\langle A,B\rangle}\hat{\sigma}^{z}_{A}\hat{\sigma}^{z}_{B}+\frac{V}{2}\sum_{A}\hat{\sigma}^{z}_{A}+\frac{\Omega}{2}\sum_{A}\hat{\sigma}^{x}_{A},$ (4)
+[…]
 Discussion and Outlook.—We have investigated the long-time mean-field dynamics of large-scale spin systems and uncovered a prethermal plateau with emergent $\mathbb{Z}_{2}$ gauge symmetry, whose breakdown follows a universal surface growth of nucleation regions characterized by non-linear growth exponents – opening a path to explore KPZ dynamics in quantum simulators Wei et al.
@@
 (2025a).
+Our observations provide a guide to develop rigorous theories of slow thermalization in LGTs Michailidis et al.
+(2018); Causer et al.
+(2024); Marić et al.
+(2026).
 While ED confirms the plateau’s stability in small systems, existing numerical methods, including DTWA, cannot capture the quantum evolution, highlighting quantum simulators Altman et al.
-(2021) – such as Rydberg tweezer arrays, trapped ions, or superconducting qubits – as promising platforms to experimentally probe these many-body dynamics.
+(2021) – such as Rydberg tweezer arrays, trapped ions, or superconducting qubits – as promising platforms to experimentally probe these many-body dynamics and to benchmark our method for simulating large-scale gauge theories.
 Acknowledgements.—We thank Rahul Nandkishore, Alexander Baumgärtner, Adam Kaufman, Christian Groß, Joanna Lis, Hannes Pichler, Mu Qiao and Ruben Verresen for fruitful discussions.